What are the steps involved in calculating how many atoms exist in 100g of lithium?


1
Expert's answer
2022-01-13T03:50:05-0500

n = m / M

M (Li) = 6.941 g/mol

n (Li) = 100 / 6.941 = 14.41 mol


1 mol contains 6.022×1023 elementary entities of the substance.

N (Li) = 14.41 mol x 6.022×1023 = 86.76 x ×1023 atoms